# Using the antenna impedance to estimate soil electrical parameters for the MIST global 21-cm experiment

### v1

@@ 5.3 Uncertainties in the best-fit soil parameters
5.4 Comparison with expectations
In this initial effort to demonstrate the antenna impedance technique we do not attempt to offer a geophysical interpretation of our results shown in Table 3. We note, however, that our results are in broad agreement with expectations from the soil composition at the test sites. Sand and gravel, found at the surface of Sites A and B, have reported conductivity $\lesssim 0.01$ Sm-1 and relative permittivity in the range 2similar-toabsent2$\sim$ 2∼ 230303030. For clay, found at the surface of Site C, the conductivity is in the range $10^{-4}$$1$ Sm-1 and the relative permittivity in the range 2similar-toabsent2$\sim$ 2∼ 240404040. Within these ranges, the specific electrical parameter values for soils composed of sand, gravel, or clay are strongly determined by the soil moisture (Reynolds 2011).
Our conductivity estimate for Site A is not expected to match very closely the results from Alegría Calvo et al. (2013). First, being $650$ m apart, the two spots do not necessarily have similar soil characteristics. Second, while our result is valid for the range 25252525125125125125 MHz, in ERT the resistivity —inverse of conductivity— is estimated in the DC regime [^3] Electrodes are used to inject direct currents into the soil and measure the resulting voltages between different spots (Slichter 1933; Koh et al. 1997; IEEE 2011; Herring et al. 2023). Sometimes, the polarization of the injected current is reversed at frequencies $\lesssim 100$ Hz to reduce the effect of electrolytic polarization (build up of anions and cations around the electrodes) and telluric currents (currents in the ground mainly produced by the presence and fluctuation of the Earth’s magnetosphere) (IEEE 2011; Reynolds 2011; Clement et al. 2020).. Despite not expecting a good agreement, such an agreement would be mathematically possible when accounting for the spatial dependence of the results. Our result represents a weighted average of the soil conductivity over space, where the weights corresponds to the sensitivity of the antenna at different coordinates in the soil. Reporting the pattern for the antenna sensitivity in the soil falls outside the scope of this paper. However, using a reasonable set of weights for the conductivities of the four layers of the ERT model it would be possible to obtain a weighted average close to our value. A realistic situation consists of the antenna having higher sensitivity toward the top of the soil. Therefore, assigning as an example weights of $96\%$ to the top layer, $4\%$ to the second layer, and $0\%$ to the other layers, produces a weighted average of $0.0076$ Sm-1, consistent with our result $0.0070\pm 0.0014$ Sm-1.
@@ 6 Discussion and summary
Our results demonstrate that the impedance of the MIST antenna can be used for the electrical characterization of the soil in the same frequency range as the sky observations. The electromagnetic performance of MIST is very sensitive to the soil electrical properties due to MIST’s strategy of observing the sky without a metal ground plane under the antenna. Therefore, an accurate characterization of the soil at the observation sites is critical for the experiment’s success. The technique demonstrated here is particularly attractive for MIST since the measurement of the antenna impedance is already included in the calibration program of the experiment (Monsalve et al. 2024a).
@@
The antenna impedance technique was used to estimate the soil electrical conductivity and relative permittivity at three test sites in the Greater Concepción area, Chile. The parameter estimates were determined by fitting impedance measurements with models from electromagnetic simulations of the antenna and soil. The soil was simulated as homogeneous. The measurements and simulations were conducted in the range 25252525125125125125 MHz. For the three sites, the electrical conductivity was found between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulated impedances have a better than $10\%$ agreement with the data relative to the peak absolute values of the resistance and reactance across our frequency range.
@@
This technique has been used in the past to characterize the environment around an antenna (e.g., Wakita et al. 2000; Lenler-Eriksen & Meincke 2004; Berthelier et al. 2005; Le Gall et al. 2006) but, to our knowledge, we are the first to use it in the context of $21$-cm cosmology. Measurements of soil electrical parameters reported so far by $21$-cm experiments have been done using other techniques (Sutinjo et al. 2015; Spinelli et al. 2022). To demonstrate the impedance technique, in this initial effort we have focused on computational aspects, in particular the electromagnetic simulations and parameter estimation. We expect future refinements in the technique to enable us to reach the soil characterization accuracy required for $21$-cm cosmology.
@@
The accuracy required in the soil parameter estimates to detect the global $21$-cm signal depends on the experiment’s design and data analysis strategy used (Pattison et al. 2024). Nonetheless, for a variety of scenarios, Spinelli et al. (2022) and Monsalve et al. (2024b) showed using signal extraction simulations that for realistic soils the accuracy required is typically better (lower) than $10\%$. The lowest parameter uncertainty reported in this paper is $20\%$ at $68\%$ probability. To characterize the soil at the sky observation sites we will use the impedance technique with the full MIST instrument instead of only the antenna panels. Using the full instrument will lead to lower uncertainties in the measurements and simulations, and therefore lower uncertainties in the soil parameters.
@@
The impedance measurements done with the full instrument are expected to be more accurate for three main reasons. First, the instrument carries out the measurements autonomously, thus eliminating the impact of the human presence and operations. Second, the impedance is measured with a better VNA and calibrated with accurately modeled OSL standards, all of which are contained in the receiver box located under the antenna (Monsalve et al. 2024a). Third, the instrument does not use external cables or the adapter required for this paper, which introduce time instabilities that cannot be tracked and corrected.
The simulations of the full instrument are also expected to be more accurate. The cable and adapter used for this paper have a complicated geometry, which introduces inaccuracies in the simulations. These inaccuracies will be avoided in the simulations of the instrument. Further, the simulation geometry will be designed using accurate measurements of the instrument’s physical dimensions done in the field, instead of relying on nominal values and laboratory measurements as in this paper.
@@
In addition to increasing the accuracy in the measurements and simulations, quantifying their uncertainties represents a priority for future implementations of this technique. Robust uncertainty estimates will enable us to select between competing models for the soil as we attempt to make the model more realistic, and determine reliable uncertainties for the soil parameters.
@@
This soil characterization technique is directly relevant for global $21$-cm experiments planning to observe the sky from the surface of the lunar farside (e.g., Bale et al. 2023). The electrical characteristics of the lunar regolith vary significantly with location (Feng et al. 2020; Siegler et al. 2020). Therefore, to reach the calibration accuracy required for $21$-cm cosmology, the conductivity and relative permittivity of the regolith will have to be determined at each observation site. Measurements of the antenna impedance primarily conducted for radiometer calibration could also be used for in-situ regolith characterization.
Acknowledgements.
@@
We thank Roberto Vives and Benjamín Fernández for their help with the antenna construction and measurements. We also thank Matheus Pessôa for his work during early tests of the antenna impedance technique. We acknowledge support from ANID Chile Fondo 2018 QUIMAL/180003, Fondo 2020 ALMA/ASTRO20-0075, Fondo 2021 QUIMAL/ASTRO21-0053, and Fondo 2022 ALMA/31220012. We acknowledge support from FONDECYT Fondo Iniciación 11221231. We acknowledge support from Universidad Católica de la Santísima Concepción Fondo UCSC BIP-106. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2019-04506, RGPNS 534549-19. We acknowledge the support of the Canadian Space Agency (CSA) [21FAMCGB15]. This research was undertaken, in part, thanks to funding from the Canada 150 Research Chairs Program. This research was enabled in part by support provided by SciNet and the Digital Research Alliance of Canada.
@@ Appendix A Geometry of the Feko simulations
Here we provide details of the Feko simulation geometry, which is shown in Fig. 4. The geometry was designed based on nominal dimensions and laboratory measurements of the antenna assembly. This assembly was replicated at the three test sites as closely as possible.
@@
In the simulated geometry, the surface of the soil is perfectly flat and aligned with the $xy$-plane at $z=0$. The antenna panels are simulated as aluminum boxes of $1.2$ m in length, $60$ cm in width, $3$ mm in thickness, and $2$ cm in horizontal separation. The length and width of the panels are aligned parallel to the $xy$-plane at a height $z=50$ cm. The long dimension of the panels is aligned with the $x$-axis. The gap between the panels is centered at $(x,y,z)=(0,0,50\;\mathrm{cm})$.
@@
The wires of the adapter in the gap between the panels are simulated as copper cylinders with a diameter of $0.64$ mm. The wires make contact with the panels while aligned with the $x$-axis at the center of the gap. As they leave the panels, the wires bend in the $+y$ direction, where they extend for $6.3$ cm before reaching the SMA connector. The connector is simulated as a gold cuboid of $1.6$ cm in length and $6$ mm in width and thickness. The electrical excitation of the antenna is applied by a voltage port located at the junction between one of the wires and the SMA connector. The location of the port mimics the impedance calibration point for the real antenna at the SMA connector.
@@
The cable is simulated as a $2$-mm diameter copper cylinder leaving the antenna in the $+y$ direction. This cylinder begins at the output of the SMA connector and runs along the middle of the gap between the panels up to the edge of the panels. At the edge of the panels, the cylinder bends downward $18.3^{\circ}$ and continues in diagonal until it reaches a lowest height of $5$ mm above the soil. This point occurs at a distance of $1.5$ m from the edge of the panels in the $+y$ direction. The cylinder then starts running parallel to the soil in the $+y$ direction at a height of $5$ mm until it reaches a total length of $20$ m.
@@
In the simulations we use the following conductivities (Pozar 2005): for aluminum, $3.816\times 10^{7}$ Sm-1; for copper, $5.813\times 10^{7}$ Sm-1; and for gold, $4.098\times 10^{7}$ Sm-1. The support structure of the antenna, made out of fiberglass and plastic, does not affect the antenna impedance and is not included in the simulations. Other features not included are the panel mounting holes. These holes would significantly increase the complexity of the geometry and computation time but according to our tests do not have a noticeable effect on the impedance.

### v2

@@ 6 Discussion and summary
Our results show that the impedance of an antenna used to conduct ground-based global $21$ cm observations can be used for the electrical characterization of the soil in the same frequency range. We demonstrate this technique using the antenna of the MIST experiment. The electromagnetic performance of MIST is very sensitive to the soil electrical properties due to MIST’s strategy of observing the sky without a metal ground plane under the antenna. Therefore, an accurate characterization of the soil at the observation sites is critical for the experiment’s success. The technique demonstrated here is particularly attractive for MIST since the measurement of the antenna impedance is already included in the calibration program of the experiment (Monsalve et al., 2024a).
@@
The antenna impedance technique was used to estimate the soil electrical conductivity and relative permittivity at three test sites in the Greater Concepción area, Chile. The parameter estimates were determined by fitting impedance measurements with models from electromagnetic simulations of the antenna and soil. The soil was simulated as homogeneous. The measurements and simulations were conducted in the range 2525125125 MHz. For the three sites, the electrical conductivity was found between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulated impedances have a better than $10\%$ agreement with the data relative to the peak absolute values of the resistance and reactance across our frequency range. Interpreting the fit residuals will require knowledge about the measurement uncertainties in future implementations of the technique.
@@
This technique has been used in the past to characterize the environment around an antenna (e.g., Wakita et al., 2000; Lenler-Eriksen & Meincke, 2004; Berthelier et al., 2005; Le Gall et al., 2006) but, to our knowledge, we are the first to use it in the context of $21$ cm cosmology. Measurements of soil electrical parameters reported so far by $21$ cm experiments have been done using other techniques (Sutinjo et al., 2015; Spinelli et al., 2022). To demonstrate the impedance technique, in this initial effort we have focused on computational aspects, in particular the electromagnetic simulations and parameter estimation. The total number of simulations performed to estimate the parameters was $\leq 602$ for each of the three test sites. Although the simulations are computationally expensive, not having to include a large metal ground plane makes the MIST simulations simpler and faster, which is a critical factor to consider when hundreds or thousands of simulations are required to characterize the soil.
The soil parameter accuracy required to detect the global $21$ cm signal depends, among other factors, on the properties of the soil itself [^5] For instance, the accuracy requirement on $\sigma$ is in general a function of $\sigma$ and $\epsilon_{r}$., the characteristics of the instrument, and the data analysis strategy used (Spinelli et al., 2022; Monsalve et al., 2024b; Pattison et al., 2025). It is therefore difficult to determine accuracy requirements that are applicable beyond a specific experimental scenario and analysis approach. In Monsalve et al. (2024b) we estimated first-order soil parameter accuracy requirements for MIST when using a simple cosmological analysis approach and assuming homogeneous and two-layer models for the soil. We found that for homogeneous soils, an accuracy of $10\%$ in $\sigma$ and $\epsilon_{r}$ was sufficient to obtain reasonably accurate estimates for the global $21$ cm signal parameters (first column of Figure $1$ in Monsalve et al. (2024b)). Two-layer soils generally increased the accuracy requirements to percent level (second through fifth columns of the same figure).
@@
The parameter uncertainties reported in this paper —in most cases better than $33\%$ at $68\%$ probability— are interesting considering that Monsalve et al. (2024b) hints at the possibility that errors $10\%$ might be tolerable for homogeneous soils. Nonetheless, acknowledging the limitations of homogeneous soil models and anticipating additional challenges in the analysis of real observations, we aim to significantly increase the accuracy and precision of our soil parameter estimates. To characterize the soil at the sky observation sites, we will use the impedance technique with the full MIST instrument instead of only the antenna panels connected with exposed cables and wires as in this paper. Using the technique with the full instrument will lead to higher accuracy in the measurements and electromagnetic simulations, and consequently higher accuracy in the soil parameters.
@@
The impedance measurements done with the full instrument are expected to be more accurate for three main reasons. First, the instrument carries out the measurements autonomously, thus eliminating the impact of the human presence and operations. Second, the impedance is measured with a better VNA and calibrated with accurately modeled OSL standards, all of which are contained in the receiver box located under the antenna (Monsalve et al., 2024a). Third, the instrument does not use external cables or the adapter required for this paper, which introduce time instabilities that cannot be tracked and corrected. The electromagnetic simulations of the full instrument are also expected to be more accurate. The cable and adapter used for this paper have a complicated geometry, which introduces inaccuracies in the simulations. These inaccuracies will be avoided in the simulations of the instrument. Further, the simulation geometry will be designed using accurate measurements of the instrument’s physical dimensions and orientation done in the field, instead of relying on nominal values and laboratory measurements as in this paper.
@@
In addition to increasing the accuracy in the measurements and simulations, quantifying their uncertainties represents a priority for future implementations of this technique. Robust uncertainty estimates for the measurements and simulations will enable us to: (1) use rigorous statistical methods to select between competing models for the soil as we attempt to make the model more realistic, and (2) determine reliable uncertainties for the soil parameters.
@@
The soil characterization technique presented in this paper is directly relevant for global $21$ cm experiments planning to observe the sky from the surface of the lunar farside (e.g., Bale et al., 2023). The electrical characteristics of the lunar regolith vary significantly with location (Feng et al., 2020; Siegler et al., 2020). Therefore, to reach the calibration accuracy required for $21$ cm cosmology, the conductivity and relative permittivity of the regolith will have to be determined at each observation site. Measurements of the antenna impedance primarily conducted for radiometer calibration could also be used for in-situ regolith characterization.
@@
We thank Roberto Vives and Benjamín Fernández for their help with the antenna construction and measurements. We also thank Matheus Pessôa for his work during early tests of the antenna impedance technique. We acknowledge support from ANID Fondo 2018 QUIMAL/180003, Fondo 2020 ALMA/ASTRO20-0075, Fondo 2021 QUIMAL/ASTRO21-0053, and Fondo 2022 ALMA/31220012. We acknowledge support from ANID FONDECYT Iniciación 11221231 and ANID FONDECYT Regular 1251819. We acknowledge support from Universidad Católica de la Santísima Concepción Fondo UCSC BIP-106. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2019-04506, RGPNS 534549-19. We acknowledge the support of the Canadian Space Agency (CSA) [21FAMCGB15]. This research was undertaken, in part, thanks to funding from the Canada 150 Research Chairs Program. This research was enabled in part by support provided by SciNet and the Digital Research Alliance of Canada.
@@ Appendix A Geometry of the Feko simulations
A.1 Simple geometry
@@
We initially conducted a set of Feko simulations with a simple geometry to evaluate how well the simulated impedance could fit the measurements without incorporating into the geometry small features that increase the computation time. The simple geometry only consists of the antenna panels, an idealized excitation port, and the soil. The antenna panels are simulated as aluminum boxes of nominal dimensions: $1.2$ m in length, $60$ cm in width, $3$ mm in thickness, and $2$ cm in horizontal separation. The length and width of the panels are aligned parallel to the $xy$-plane at a height $z=50$ cm. The long dimension of the panels is aligned with the $x$-axis. The gap between the panels is centered at $(x,y,z)=(0,0,50\;\mathrm{cm})$. The panels are connected by an ideal $2$ cm-long wire centered at $(x,y,z)=(0,0,50\;\mathrm{cm})$ and aligned with the $x$-axis. A voltage port representing the electrical excitation of the antenna is placed in the middle of the wire. The simulated soil is homogeneous, and its surface is perfectly flat and aligned with the $xy$-plane at $z=0$.
Figure 8: Comparison between the impedance measurements and Feko simulations done with (left) a simple geometry and (right) a full geometry of the MIST antenna and measurement setup. The soil parameter combinations used in these simulations span a wide realistic space. The simple simulation geometry —which only includes the antenna panels, an idealized excitation port, and the soil— leads to large systematic differences with the data. The full simulation geometry —which also includes the wires and SMA connector in the gap between the panels, as well as the $20$ m cable— improves the overall agreement.
Panels (a)–(d) of Figure 8 show four simulations conducted with the simple geometry, as well as the measurements at the three test sites for reference. In the simulations, the soil was assigned $(\sigma,{\epsilon}_{r})=(0.001\;\mathrm{Sm}^{-1},1)$, $(0.001\;\mathrm{Sm}^{-1},20)$, $(0.1\;\mathrm{Sm}^{-1},1)$, and $(0.1\;\mathrm{Sm}^{-1},20)$. These combinations span a wide realistic parameter space. The panels in the figure show that across some frequency ranges the simulations with the simple geometry differ from the measurements by amounts that cannot be attributed to the soil parameters. The clearest example is in $X$, where above $85$ MHz the simulations and measurements converge to different values, which at $125$ MHz are separated by $\sim$$35$ $\Omega$. Systematic differences between simulations and measurements can also be seen in: (1) $|\Gamma|$ at $\sim$8585120120 MHz; (2) $\angle\Gamma$ above $\sim$$85$ MHz; (3) $R$ at $\sim$8080105105 MHz; and (d) $X$ at $\sim$25254545 MHz.
The systematic differences observed between the simulations with the simple geometry and the data motivated us to make the geometry more realistic. The improved geometry, which we here call ‘full’, is described in Appendix A.2. Panels (e)–(h) of Figure 8 show simulations with the full geometry. In these simulations, the soil parameters take the same values as in the simulations with the simple geometry shown in panels (a)–(d). The simulations with the full geometry agree better with the measurements, and some remaining systematic differences are significantly smaller than when using the simple geometry. These results made us choose the full geometry for the simulations in this paper.
A.2 Full geometry
In the full simulation geometry, the antenna panels and soil have the same characteristics as in the simple geometry. However, we remove the idealized excitation port from the gap between the panels. We then incorporate the wires and SMA connector of the adapter between the panels, as well as the $20$ m coaxial cable.
@@ A.2 Full geometry
The wires of the adapter are simulated as copper cylinders with a diameter of $0.64$ mm. The wires make contact with the panels while aligned with the $x$-axis at the center of the gap. As they leave the panels, the wires bend in the +y+y direction, where they extend for $6.3$ cm before reaching the SMA connector. The connector is simulated as a gold cuboid of $1.6$ cm in length and $6$ mm in width and thickness. The electrical excitation of the antenna is applied by a voltage port located at the junction between one of the wires and the SMA connector. The location of the port mimics the impedance calibration point for the real antenna at the SMA connector.
@@
The cable is simulated as a $2$ mm diameter copper cylinder leaving the antenna in the +y+y direction. This cylinder begins at the output of the SMA connector and runs along the middle of the gap between the panels up to the edge of the panels. At the edge of the panels, the cylinder bends downward $18.3^{\circ}$ and continues in diagonal until it reaches a lowest height of $5$ mm above the soil. This point occurs at a distance of $1.5$ m from the edge of the panels in the +y+y direction. The cylinder then starts running parallel to the soil in the +y+y direction at a height of $5$ mm until it reaches a total length of $20$ m.
@@
In the simulations we use the following conductivities (Pozar, 2005): for aluminum, $3.816\times 10^{7}$ Sm-1; for copper, $5.813\times 10^{7}$ Sm-1; and for gold, $4.098\times 10^{7}$ Sm-1. The support structure of the antenna, made out of fiberglass and plastic, does not affect the antenna impedance and is not included in the simulations. Other features not included are the mounting holes of the antenna panels. These holes would significantly increase the complexity of the geometry and computation time but according to our tests do not have a noticeable effect on the impedance.
Appendix B Estimating the soil parameters and their uncertainties
To determine the best-fit soil parameter values we use the simplicial homology global optimization (SHGO) algorithm (Endres et al., 2018; Virtanen et al., 2020). SHGO enables derivative-free, constrained and unconstrained optimization. SHGO minimizes the cost function using global and local searches. For the global search, SHGO incorporates the simplicial (Hatcher, 2002), Sobol (Sobol, 1967; Joe & Kuo, 2008), and Halton (Halton, 1960) sampling methods, and can also accept externally provided sampling points. In this work we use the Sobol and Halton low-discrepancy sequences for the global search. These sequences are deterministic sets of sampling points with mathematical properties comparable to, and in some cases better than, the points from the popular Latin hypercube sampling (McKay et al., 1979; Renardy et al., 2021). For the local search, we use sequential least squares quadratic programming (Kraft, 1988; Nocedal & Wright, 2006), which is the default method in SHGO.
Figure 9: Samples produced in the initial fitting step. The colors of the samples represent the SSD. The three columns correspond to the three test sites. The top (bottom) row corresponds to the Sobol (Halton) samples. In each panel, the sample with the lowest SSD, denoted by SSD, is highlighted using a red circle. In all cases, SSD occurs for $\sigma<0.06$ Sm-1.
Figure 10: Same as Fig. 9 but for the final fitting step. This step enables the identification of the best-fit values with higher accuracy than the initial step. The color scales represent different SSD ranges for each of the three sites.
During the parameter fits we use the default settings that define the stopping criterion in SHGO. With this choice it is necessary to fit the parameters in two steps in order to explore a wide parameter space and converge to accurate best-fit values. In the initial step we run SHGO within the wide space defined by the following parameter ranges: 011 Sm-1 for $\sigma$ and 112020 for $\epsilon_{r}$. The best-fit values found in this step are not very accurate but indicate the regions where the true best-fit values lie. In the final fitting step we run SHGO again but within smaller parameter spaces, which we select based on the best-fit values found in the initial step. To quantify the accuracy of the best-fit values we carry out the two-step fitting process twice, using separately the Sobol and Halton sampling points for the global exploration. Although conducting the fits twice roughly doubles the number of Feko evaluations, the total number is expected to be less than 1000 per site, which is lower than typical numbers required for MCMC or nested sampling.
The fit results are shown in Figures 9 and 10, and summarized in Table 4. The table lists: (1) the best-fit soil parameters, $\hat{\sigma}$ and $\hat{\epsilon}_{r}$; (2) the sum of square deviations (SSD) at the best-fit values, SSD; (3) the root mean square deviation (RMSD) between the best-fit impedance models and the data, $\mathrm{RMSD}^{\star}=\sqrt{\mathrm{SSD}^{\star}/N}$, where $N=22$; and (4) the number of evaluations, $N_{\mathrm{ev}}$. These results are presented for the initial and final fitting steps, and for the Sobol and Halton samples.
Figure 9 shows the samples and SSDs produced in the initial fitting step. The SSDs go from under 2000 $\Omega^{2}$ to over 7000 $\Omega^{2}$. For the three test sites, the SSD decreases toward low $\sigma$, while across $\epsilon_{r}$ it remains more stable. The samples that produce the SSD are indicated using red circles. For Site A, as shown in Table 4, the best-fit values found in the initial step are $(\hat{\sigma},\hat{\epsilon}_{r})=(10^{-5}\;\mathrm{Sm}^{-1},1)$ for Sobol and $(0.0054\;\mathrm{Sm}^{-1},6.95)$ for Halton. These values are used as reference to determine adequate parameter ranges for the final fitting step. The ranges selected for the final step are 00.020.02 Sm-1 for $\sigma$ and 111010 for $\epsilon_{r}$, which contain the initial best-fit values. For Sites B and C we follow a similar approach but decide to only restrict the $\sigma$ range. For these sites, the best-fit values found for $\epsilon_{r}$ in the initial step are higher than for Site A and the variation of the SSD across $\epsilon_{r}$ is low. We thus prefer to maintain 112020 as the range for $\epsilon_{r}$ in the final step. The $\sigma$ range is restricted to 00.040.04 Sm-1 for Site B and 00.080.08 Sm-1 for Site C.
Table 4: Results of the parameter estimation with SHGO.
Initial fitting step Final fitting step
Site Sampling $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD $\Omega^{2}$] RMSD $\Omega$] $N_{\mathrm{ev}}$ $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD $\Omega^{2}$] RMSD $\Omega$] $N_{\mathrm{ev}}$
A Sobol $0^{-5} 11$ $1132.0$ $7.2$ $131$ $0.0070$ $1.63$ $513.3$ $4.8$ $183$
A Halton $0.0054$ $6.95$ $1880.6$ $9.2$ $103$ $0.0064$ $1.82$ $531.8$ $4.9$ $151$
B Sobol $0.0234$ $8.57$ $1869.1$ $9.2$ $131$ $0.0144$ $9.61$ $1822.6$ $9.1$ $220$
B Halton $0.0210$ $8.12$ $1861.1$ $9.2$ $103$ $0.0165$ $9.77$ $1815.5$ $9.1$ $148$
C Sobol $0.0547$ $18.66$ $930.0$ $6.5$ $137$ $0.0494$ $12.73$ $825.9$ $6.1$ $188$
C Halton $0.0523$ $16.33$ $864.2$ $6.3$ $106$ $0.0479$ $14.69$ $834.1$ $6.2$ $148$
Figure 10 shows the samples and SSDs produced in the final fitting step. The final parameter spaces selected based on the initial fitting step successfully capture the regions of lowest SSD and enable a denser sampling of these regions. The denser sampling leads to a higher accuracy in the best-fit values, which is reflected in the agreement at the $\sim$$15\%$ level or better between the Sobol and Halton best-fit values. We report as the final best-fit values those that produce the lowest SSD considering both the Sobol and Halton sample sets. For Sites A, B, and C, the final best-fit values are $(\hat{\sigma},\hat{\epsilon}_{r})=(0.0070\;\mathrm{Sm}^{-1},1.63)$, $(0.0165\;\mathrm{Sm}^{-1},9.77)$, and $(0.0494\;\mathrm{Sm}^{-1},12.73)$, respectively. The corresponding SSD are $513.3$, $1815.5$, and $825.9$ $\Omega^{2}$. The total number of sampled points considering the two fitting steps and the Sobol and Halton samples is $\leq 602$ for each of the three sites.